Calculate the equilibrium level level of income if c=100 million +0.8y if ī=125 million
1
Expert's answer
2019-07-29T08:11:48-0400

IF:

c=100 million +0.8Y and I=125 million


Y=C+I=100+0.8Y+125=0.8Y+225

0.2Y=225

Y=225/0.2=1125